Circular No. 7120

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1819–254, XTE J1743–363, XTE J1710–281, XTE J1723–376 C. B. Markwardt, National Research Council and Goddard Space

Flight Center (GSFC); and J. H. Swank and F. E. Marshall, GSFC, report the detection of two new x-ray transients: “XTE J1819–254, whose position is consistent with that of the independently discovered SAX J1819.3–2525 (cf. IAUC 7119), had fluxes of 20, 37, 3, and 24 mCrab in the band 2–10 keV on Feb. 18.78, 22.65, 25.84, and Mar. 1.44 UT, respectively, and < 1 mCrab before that time. XTE J1743–363 has been active since the beginning of Feb. with intensities ranging from 3 to 15 mCrab. Variability of as much as 25 percent (r.m.s.) on approximately 1-min timescales has been seen. The best-fit position is α = 17^{h}42^{m}58^{s}, δ = −36^{o}19′ (equinox 2000.0), with an estimated uncertainty of 10′. Recent transients XTE J1710–281 (IAUC 6998) and XTE J1723–376 (IAUC 7103) remain active, with XTE J1710–281 showing evidence of an occultation (duration < 800 s). Semiweekly RXTE monitoring observations of the Galactic-center region will continue.”

COMET P/1998 U3 (JÄGER)

Ephemeris extension to IAUC 7062 (elements MPC 33650):

1999 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 3 6^{h}29.^{m}56 +24^{o}10.^{′}6 1.521 2.135 114.^{o}8 24.^{o}9 10.7
13 6 39.25 +22 13.9 1.619 2.134 107.1 26.4 10.8
23 6 51.20 +20 25.2 1.725 2.137 100.1 27.3 11.0
Apr. 2 7 04.96 +18 42.0 1.837 2.145  93.6 27.7 11.1
12 7 20.10 +17 02.0 1.953 2.157  87.5 27.7 11.3
22 7 36.29 +15 22.9 2.073 2.172  81.9 27.3 11.5
May 2 7 53.23 +13 43.1 2.194 2.192  76.6 26.6 11.6
12 8 10.69 +12 01.2 2.317 2.215  71.5 25.6 11.8
22 8 28.48 +10 16.6 2.441 2.242  66.7 24.5 11.9
June 1 8 46.45 + 8 28.7 2.564 2.272  62.0 23.2 12.1
11 9 04.47 + 6 37.5 2.687 2.305  57.4 21.8 12.3
21 9 22.48 + 4 42.9 2.809 2.341  53.0 20.3 12.4
July 1 9 40.38 + 2 45.4 2.928 2.380  48.5 18.7 12.6
11 9 58.14 + 0 45.3 3.045 2.422  44.2 17.0 12.8
21 10 15.73 − 1 17.0 3.158 2.465  39.8 15.3 12.9
31 10 33.13 − 3 20.8 3.267 2.511  35.5 13.6 13.1
Aug. 10 10 50.32 − 5 25.6 3.371 2.559  31.2 11.9 13.2
1999 March 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green